Instruction: Find the solution set of the following using 3 methods. Show your solutions, then submit your answer to your eLMS dropbox.


1.2x + y = 15

x - y = 9


2. 3x - 2y = 12

  2x + 5y = 72. 3x - 2y = 12

  2x + 5y = 7


1
Expert's answer
2022-06-15T14:12:02-0400

There are three ways to solve systems of linear equations in two variables:

graphing.

substitution method.

elimination method.

1.


"2x + y = 15""x - y = 9"


a.

"(8, -1)"


b.


"2x + y = 15""x - y = 9"




"2x + x-9 = 15""y = x-9"




"3x = 24""y= x-9"




"x=8""y=-1"

"(8, -1)"


c.


"2x + y = 15""x - y = 9"




"3x - 2y = 12""2x + 5y = 72"




"2x + y+x-y= 15+9""x-y=9"




"3x = 24""x-y=9"




"x=8""8-y=9"


"x=8""y=-1"

"(8, -1)"



2.


"3x - 2y = 12""2x + 5y = 72"


a.

"(\dfrac{204}{19}, \dfrac{192}{19})"


b.


"3x - 2y = 12""2x + 5y = 72"




"y = \dfrac{3}{2}x-6""2x+5(\dfrac{3}{2}x-6)=72"




"y = \dfrac{3}{2}x-6""\dfrac{19}{2}x=102"




"x=\dfrac{204}{19}""y=\dfrac{3}{2}(\dfrac{204}{19})-6"


"x=\dfrac{204}{19}""y=\dfrac{192}{19}"


"(\dfrac{204}{19}, \dfrac{192}{19})"


c.


"3x - 2y = 12""2x + 5y = 72"




"6x - 4y = 24""6x + 15y = 216"




"6x+15y-(6x-4y)=216-24""3x-2y=12"




"19y=192""x=\dfrac{2}{3}y+4"




"x=\dfrac{204}{19}""y=\dfrac{192}{19}"

3.


"3x - 2y = 12""2x + 5y = 7"


a.


"(\dfrac{74}{19}, -\dfrac{3}{19})"


b.


"3x - 2y = 12""2x + 5y = 7"




"y = \dfrac{3}{2}x-6""2x+5(\dfrac{3}{2}x-6)=7"




"y = \dfrac{3}{2}x-6""\dfrac{19}{2}x=37"




"x=\dfrac{74}{19}""y=\dfrac{3}{2}(\dfrac{74}{19})-6"


"x=\dfrac{74}{19}""y=-\dfrac{3}{19}"


"(\dfrac{74}{19}, -\dfrac{3}{19})"


c.


"3x - 2y = 12""2x + 5y = 7"




"6x - 4y = 24""6x + 15y = 21"




"6x+15y-(6x-4y)=21-24""3x-2y=12"




"19y=-3""x=\dfrac{2}{3}y+4"




"x=\dfrac{74}{19}""y=-\dfrac{3}{19}"

"(\dfrac{74}{19}, -\dfrac{3}{19})"